### v1

@@ 4.1 LiDAR-like experiment
Table 1 shows the results of testing the networks trained with the LiDAR-like dataset with the EVO dataset. The confusion matrix of each network is presented in Table 2. It can be seen that, regardless of having an overall good accuracy, the networks still struggle to differentiate understorey from terrain, especially when the vegetation is on a near-ground level. Despite this, we can notice that the networks perform considerably well in segmenting the trees from the rest of the categories and even segmenting trunks from the canopy. PointNeXt had a better result in classifying the points in the EVO dataset and segmenting each category, and therefore it seems more suitable for forest environments.
@@ 4.2 Camera-like experiment
Table 3 shows the results of testing the networks trained with the Camera-like dataset with the EVO dataset with occlusion. The confusion matrices of each network can be seen in Table 4. Similar to the previous case, the networks struggle to differentiate terrain points from understorey points, and as few trunk points remain visible, especially the ones at the base of the trunk, it is also often confused with understorey points. We can see that the overall performance is significantly lower than in the case without occlusion. PointMAE achieves slightly better accuracy than the other networks, although all have similar responses.
@@ 5 Conclusions and future work
In this work, we developed an open-source simulator based on Unity Engine that generates realistic forest scenes procedurally. With it, we have created synthetic point-based datasets, with each point labeled into one of the categories: terrain, trunk, canopy, and understorey (including grass, bushes, and other vegetation that are not trees). We then employed these datasets to train four state-of-the-art deep-learning point-based networks. Finally, we tested and compared them in the real forest EVO dataset. The results show that synthetic data can be used to train deep-learning networks for posterior forest segmentation with real data. Among the tested networks, PointNeXt seemed to give better results when trained with the LiDAR-like dataset than the other networks, whereas PointMAE obtained slightly better accuracy when trained with the Camera-Like dataset.
As future work, we will use synthetic data to pre-train the deep learning networks and then fine-tune them using real data, expecting better accuracy. This result is also relevant since it would allow data collection in only a portion of the forest area of interest to conduct deep learning network training.

### v2

@@ 4.1 LiDAR-like experiment
Table 1 shows the results of testing the networks trained with the LiDAR-like dataset with the Evo dataset. The confusion matrix of each network is presented in Table 2. It can be seen that, regardless of having an overall good accuracy, the networks still struggle to differentiate understorey from terrain, especially when the vegetation is on a near-ground level. Despite this, in Section 4.3 we show that the networks perform well in segmenting the trees from the rest of the categories. PointNeXt had a better result in classifying the points in the Evo dataset and segmenting each category, and therefore it seems more suitable for LiDAR datasets.
@@ 4.2 Camera-like experiment
Table 3 shows the results of testing the networks trained with the Camera-like dataset with the Evo dataset with occlusion. The confusion matrices of each network can be seen in Table 4. Similar to the previous case, the networks struggle to differentiate terrain points from understorey points, and as few trunk points remain visible, especially the ones at the base of the trunk, it is also often confused with understorey points. We can see that the overall performance is significantly lower than in the case without occlusion. Again, when considering only two categories, Tree and Non-tree, the overall accuracy achieves better results (see Section 4.3). PointMAE obtains slightly better accuracy than the other networks, although all have similar responses.
@@
4.3 Tree and Non-Tree segmentation
Table 5 shows the results obtained considering only Tree and Non-Tree categories, the first one including trunk and canopy points, and the latter including terrain and understorey points, training with both datasets, LiDAR-like and Camera-like, and testing with the Evo Dataset. It can be seen that the tree segmentation obtains a high accuracy percentage with all the considered networks, being PointBERT the one with better results.
Table 5: Results obtained considering the categories Tree and Non-Tree, testing in the Evo dataset and training in both the LiDAR-like and the Camera-like datasets.
Network LiDAR-like Dataset Camera-like Dataset
PointNeXt 0.9051 0.9330 0.8035 0.9414 0.9180 0.8765
PointBERT 0.9328 0.9449 0.8652 0.9487 0.9108 0.8773
PointMAE 0.9276 0.9416 0.8561 0.9408 0.8982 0.8596
PointGPT 0.8454 0.8797 0.7252 0.9372 0.8906 0.8497
@@ 5 Conclusions and future work
In this work, we developed an open-source simulator based on Unity Engine that generates realistic forest scenes procedurally. With it, we have created synthetic point-based datasets, with each point labeled into one of the categories: terrain, trunk, canopy, and understorey (including grass, bushes, and other vegetation that are not trees). We then employed these datasets to train four state-of-the-art deep-learning point-based networks. Finally, we tested and compared them in the real forest Evo dataset. The results show that synthetic point cloud data can be used to train deep-learning networks for posterior forest segmentation with real data. Among the tested networks, PointNeXt seemed to give better results when trained with the LiDAR-like dataset than the other networks, whereas PointMAE obtained slightly better accuracy when trained with the Camera-Like dataset. When considering only two categories, Tree and Non-tree, a higher accuracy percentage is obtained, and PointBERT achieved better results than the other networks.
As future work, we will use synthetic data to pre-train the deep learning networks and then fine-tune them using real data, expecting better accuracy, specially in the experiments that considers the four categories (terrain, trunks, canopy and understorey). This result is also relevant since it would allow data collection in only a portion of the forest area of interest to conduct deep learning network training.